OpenStudy (anonymous):

Find all solutions in the interval [0,360°) tan2θ−√3tanθ=0

OpenStudy (dumbcow):

tan 2x = sqrt3 tan x **tan 2x = 2 tan x / 1-tan(x)^2 2 tanx = sqrt3 tanx(1-tanx^2) sqrt3 tanx^3 +(2-sqrt3)tanx = 0 tanx (sqrt3 tanx^2 +(2-sqrt3)) =0 tanx = 0 --> x =0,180 sqrt3 tanx^2 = -(2-sqrt3) tanx = sqrt((sqrt3 -2)/sqrt3) = sqrt(1 - (2/sqrt3)) 1 - 2/sqrt3 < 0 x has no real solutions